Circular No. 7071

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1998 VS_{24} (LINEAR) On further examining the Nov. 10–11 observations, published on MPS

3154, of an apparently unremarkable asteroidal discovery from the LINEAR program, G. V. Williams made identifications with isolated sets of observations from Višnjan on Oct. 24, 28 (the latter having been published on MPS 2894 as one of two objects that were both erroneously identified with 1998 UD_{19}) and Nov. 26. The resulting orbit was very cometary in form. Furthermore, he recognized that the object had made an extremely close approach (< 0.01 AU) to Jupiter in Oct. 1971, with additional approaches to 0.5–0.7 AU in both 1983 and 1995. Here are representative observations:

1998 UT α_{2000} δ_{2000} m_{2} Observer
Oct. 24.84196 3^{h}03^{m}16.^{s}98 +10^{o}45′25′′.2  Višnjan
28.91355 3 00 54.41 +10 30 48.8 17.9  _{”}
Nov. 10.25724 2 53 13.55 + 9 48 34.3 18.5 LINEAR
11.28224 2 52 34.86 + 9 45 20.5 18.5  _{”}
26.73186 2 43 34.79 + 9 05 16.9  Višnjan

Following placement of the object in The NEO Confirmation Page yesterday, observations during Dec. 18.9–19.3 UT have been reported by J. Tichá and M. Tichý (Kleť), P. Pravec (Ondřejov), W. B. Offutt (Cloudcroft) and G. C. L. Aikman (Victoria), most of whom remark that the object indeed seems to be of cometary appearance. Offutt, using a 0.6-m f/7 Ritchey-Chrétien, specifically notes: “Faint, somewhat asymmetrical coma extending ∼ 6′′ north, 9′′ east, 14′′ south and 9′′ west of the central condensation, which is itself featureless, having a ‘diameter’ of ∼ 5′′ (two to three times the extent of nearby star images having a similar peak intensity) and noticeable elongation in the north-south direction.” Further details on the astrometry and orbit computation are given on MPEC 1998-Y17.

Epoch = 1998 Nov. 3.0 TT

T = 1998 Nov. 6.3700 TT  ω = 245.^{o}1677
e = 0.244762  Ω = 159.1699 }2000.0
q = 3.402590 AU  i = 5.0282

a = 4.505324 AU n^{o} = 0.1030659 P = 9.563 years

1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Dec. 13 2^{h}37.^{m}14 + 8^{o}47.^{′}4 2.606 3.407 138.^{o}5 11.^{o}0 18.4
23 2 35.43 + 8 50.8 2.713 3.409 128.1 13.1 18.6
Jan. 2 2 35.63 + 9 05.0 2.837 3.413 118.1 14.7 18.8
12 2 37.72 + 9 28.9 2.973 3.416 108.6 15.8 18.9
1998 December 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden